### v2

# Airy Beamforming for Radiative Near-Field MU-XL-MIMO: Overcoming Half-Space Blockage

Abstract
The move to next-generation wireless communications with extremely large-scale antenna arrays (ELAAs) brings the communications into the radiative near-field (RNF) region, where distance-aware focusing is feasible. However, high-frequency RNF links are highly vulnerable to blockage in indoor environments dominated by half-space obstacles (walls, corners) that create knife-edge shadows. Conventional near-field focused beams offer high gain in line-of-sight (LoS) scenarios but suffer from severe energy truncation and effective-rank collapse in shadowed regions, often necessitating the deployment of auxiliary hardware such as Reconfigurable Intelligent Surfaces (RIS) to restore connectivity. We propose a beamforming strategy that exploits the auto-bending property of Airy beams to mitigate half-space blockage without additional hardware. The Airy beam is designed to “ride” the diffraction edge, accelerating its main lobe into the shadow to restore connectivity. Our contributions are threefold: (i) a Green’s function-based RNF multi-user channel model that analytically reveals singular-value collapse behind knife-edge obstacles; (ii) an Airy analog beamforming scheme that optimizes the bending trajectory to recover the effective channel rank; and (iii) an Airy null-steering method that aligns oscillatory nulls with bright-region users to suppress interference in mixed shadow/bright scenarios. Simulations show that the proposed edge-riding Airy strategy achieves a Signal-to-Noise Ratio (SNR) improvement of over 20 dB and restores full-rank connectivity in shadowed links compared to conventional RNF focusing, virtually eliminating outage in geometric shadows and increasing multi-user spectral efficiency by approximately 35% under typical indoor ELAA configurations. These results demonstrate robust RNF multi-user access in half-space blockage scenarios without relying on RIS.
I Introduction
THE advent of 6G wireless systems is characterized by the deployment of Extremely Large-Scale Antenna Arrays (ELAA) operating at millimeter-wave (mmWave) and Terahertz (THz) frequencies [1, 2, 3]. These massive apertures extend the Radiative Near-Field (RNF) to hundreds of meters, unlocking the “distance dimension” for precise beam focusing [4, 5].
However, high-frequency RNF links are intrinsically fragile. As visually summarized in Fig. 1, ubiquitous half-space obstacles (e.g., walls) create extensive geometric shadow regions. In these scenarios, traditional near-field beams, which propagate linearly [6, 7, 8], are severely truncated by the obstacle edge. This results in a catastrophic power loss and a “singular value collapse” of the multi-user channel [9], often rendering Zero-Forcing precoding ineffective and causing service outage.
To mitigate such blockage, Reconfigurable Intelligent Surfaces (RIS) are often employed to establish virtual Line-of-Sight (LoS) links [11, 12, 13, 14]. Yet, ubiquitous RIS deployment incurs significant hardware costs, power consumption, and channel estimation overhead [15]. This motivates a critical paradigm shift towards Wavefront Engineering—shaping the electromagnetic field structure at the transmitter to navigate complex environments without relying on auxiliary relays.
Figure 1: Conceptual illustration of the proposed Airy beamforming strategy versus traditional near-field focusing in a radiative near-field (RNF) half-space blockage scenario. The traditional focused beam (red trajectory) propagates linearly and is severely truncated by the knife-edge obstacle, resulting in a communication outage for the shadowed user (UE-1). In contrast, the proposed Airy beam (blue trajectory) exploits its unique auto-bending property to ”ride” the diffraction edge, curving into the geometric shadow zone to re-establish a reliable link for UE-1, while simultaneously serving the line-of-sight user (UE-k).
Recent years have witnessed a surge of interest in Non-Diffracting (ND) beams for this purpose. The seminal work by Reddy and Jornet et al. [16] demonstrated the potential of THz Bessel beams to “self-heal” after encountering partial blockage, triggering a wave of research into structured light for wireless communications. Subsequent studies have expanded this frontier: Miao et al. [19] realized extended non-diffractive ranges using dielectric metalenses, while Yang et al. [26] explored flexible antenna arrays to dynamically shape radiation patterns. In our recent work [18], we provided a quantitative roadmap for the engineering of linear ND beams (specifically Bessel and Mathieu beams) using planar phased arrays, establishing the fundamental link-level gain regimes for these structured wavefronts [18]. Complementing these theoretical studies, Bodet et al. [30] provided essential sub-terahertz near-field channel measurements, verifying the propagation characteristics of diffractive beams.
However, a critical limitation remains: Bessel and Mathieu beams, while robust to partial occlusion, still propagate along linear trajectories. They cannot reach deep into the geometric shadow of large, optically opaque obstacles. Subsequently, the potential of Airy beams has been rapidly explored across various domains: Darsena et al. [22] analyzed their fundamental performance limits in near-field channels; Lee et al. [21] experimentally validated high-speed transmission using 3D-printed lenses; and Zhang et al. [24] proposed convex optimization frameworks for bending beam generation. In specific application scenarios, Zhao et al. [25] evaluated the efficacy of Airy beams for Terahertz wireless data centers, while Hu et al. [29] demonstrated robust underwater optical communications using RGB Airy transmitters. Beyond standard communications, the unique wavefronts of Airy beams have also been exploited for physical layer security via wavefront hopping [20], complex spatiotemporal wavepacket dynamics [28], and high-resolution phase imaging through scattering media [23, 27].
Despite this flurry of activity, a critical gap persists in the literature. Most existing studies focus on single-user link budget enhancement [17, 21] or specific non-communication applications [20, 23]. The dynamics of Multi-User (MU) interference when using these curved, diffractive beams remain largely unexplored. Unlike the linear Bessel beams discussed in [18], Airy beams possess extended asymmetric sidelobes (the “tails” essential for self-bending), which pose a severe risk of co-channel interference in MU-MIMO setups. Standard linear precoding cannot easily resolve this conflict without sacrificing the bending gain.
In this paper, we propose a physics-aware Airy beamforming framework to bridge this gap. Our main contributions are:
  • RNF Blockage Modeling: We derive a Green’s function-based model revealing that while traditional focusing fails behind knife-edge obstacles, Airy beams maintain a viable channel rank.
  • Edge-Riding Beamforming: We design an analog beamforming strategy that optimizes the Airy beam’s acceleration to graze the obstacle edge, establishing a robust sum-rate floor in the shadow.
  • Null-Steering for Mixed Scenarios: For scenarios with both shadowed and LoS users, we propose a method to align the Airy beam’s inherent nulls with the LoS user, achieving analog-domain interference suppression.
The remainder of this paper is organized as follows. Section II establishes the RNF multi-user system model based on Green’s function and introduces the Zero-Forcing precoding framework. Section III analyzes the baseline performance of traditional near-field focusing in blockage-free scenarios, highlighting the distance-domain resolution of RNF. Section IV investigates the critical impact of half-space blockage on channel singular values and demonstrates the rank restoration capability of the proposed edge-riding Airy beamforming. Section V further explores a challenging mixed blockage scenario, proposing the Airy-null-steering strategy for interference mitigation between shadowed and line-of-sight users. Finally, Section VI concludes the paper and discusses future research directions.
II System Model and Channel Characterization
We consider a downlink radiative near-field (RNF) multi-user MIMO (MU-MIMO) system operating in the millimeter-wave (mmWave) band. The Base Station (BS) employs a large-scale Uniform Linear Array (ULA) to serve $K$ single-antenna users simultaneously. The propagation environment is characterized by a half-space blockage geometry, where a knife-edge obstacle potentially obstructs the Line-of-Sight (LoS) paths for specific users.
II-A Geometry and Signal Configuration
We define the system in a two-dimensional (2D) Cartesian coordinate system $(x,z)$, where the $z$-axis denotes the longitudinal propagation direction and the $x$-axis represents the transverse dimension. The BS is equipped with an $N$-element ULA located at $z=0$, spanning an aperture $D$ with elements positioned at $x_{n}=(n-\frac{N+1}{2})d$, where $d\approx\lambda/2$ is the element spacing. The $K$ users are distributed in the RNF region (Fresnel zone, $z<2D^{2}/\lambda$) at coordinates $\mathbf{p}_{k}=(x_{k},z_{k})$.
The complex baseband signal received by the $K$ users, $\mathbf{y}\in\mathbb{C}^{K\times 1}$, is given by:
$\mathbf{y}=\mathbf{H}_{\text{phys}}\mathbf{W}_{\text{RF}}\mathbf{W}_{\text{BB}}\mathbf{s}+\mathbf{n},$ (1)
where $\mathbf{H}_{\text{phys}}\in\mathbb{C}^{K\times N}$ is the physical channel matrix, $\mathbf{W}_{\text{RF}}$ and $\mathbf{W}_{\text{BB}}$ are the analog and digital precoders, respectively. $\mathbf{s}\in\mathbb{C}^{K\times 1}$ contains the normalized data symbols ($\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}_{K}$), and $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{K})$ is the additive white Gaussian noise.
II-B Physical Channel Construction: Methodological Consistency
Accurate modeling of RNF communications requires a rigorous wave-theoretic framework. In this work, we employ two complementary modeling approaches and explicitly establish their physical consistency.
II-B1 Baseline Green’s Function Model (Blockage-Free)
For LoS scenarios without obstacles, the channel response from the $n$-th antenna element to the $k$-th user is described by the free-space Green’s function (cylindrical wave approximation):
$h_{k,n}=\frac{\lambda}{4\pi r_{k,n}}e^{-jk_{0}r_{k,n}},$ (2)
where $k_{0}=2\pi/\lambda$ is the wavenumber and $r_{k,n}=\|\mathbf{p}_{k}-(x_{n},0)\|_{2}$. This captures the precise spherical wavefront curvature in the near field.
II-B2 Fresnel Diffraction Integration (With Blockage)
To account for half-space blockage and the auto-bending propagation of Airy beams, we adopt a wave-optics approach based on the Fresnel diffraction integral. The field at the receiver depth $z_{k}$ is computed by propagating the aperture field through the obstacle plane:
$E(x,z_{k})=\mathcal{P}_{\Delta z_{2}}\{\mathcal{M}(x)\cdot\mathcal{P}_{\Delta z_{1}}\{E_{0}(x)\}\},$ (3)
where $\mathcal{P}_{z}$ is the Fresnel propagation operator, and $\mathcal{M}(x)$ is the transmittance mask (1 for free space, 0 for obstacle). Here, ${\Delta z_{1}}=z_{\text{obs}}$ and ${\Delta z_{2}}=z_{k}-z_{\text{obs}}$ denote the propagation distances before and after the obstacle plane, respectively.
Remark 1 (Consistency of Modeling): It is well-established that in the paraxial regime ($z\gg\lambda$), the discrete summation of Green’s functions converges to the continuous Fresnel diffraction integral. Therefore, the use of Green’s functions for analytical insights (Sec. III) and Fresnel diffraction for blockage simulations (Secs. IV and V) ensures physical consistency throughout the paper.
II-C Hybrid Beamforming Strategy
II-C1 Analog Beamforming: Codebook and Optimization
The analog precoder $\mathbf{W}_{\text{RF}}=[\mathbf{w}_{1},\dots,\mathbf{w}_{K}]\in\mathbb{C}^{N\times K}$ maps each RF chain to a specific wave mode. We focus on two strategies:
  • Traditional Near-Field Focusing: Matches the spherical phase of the user’s location via phase conjugation: $w_{\text{trad}}^{(n)}=\frac{1}{\sqrt{N}}e^{jk_{0}r_{k,n}}$.
  • Proposed Airy Beamforming: We impart a cubic phase modulation derived from the paraxial finite-energy Airy beam formulation [10]:
    $\phi_{\text{Airy}}(x_{n})=\frac{k_{0}}{2F}x_{n}^{2}-k_{0}\sin(\theta)x_{n}+\frac{2\pi}{3\lambda}B\left(\frac{x_{n}}{F}\right)^{3}.$ (4)
    Here, $\theta$ controls the launch angle, $F$ is a virtual focal parameter, and $B$ determines the transverse acceleration (bending). Optimization Insight: Unlike static parameter settings, the tuple $(B,F,\theta)$ provides degrees of freedom to manipulate the beam trajectory and side-lobe positions. In multi-user scenarios (Sec. V), these parameters can be tuned to align the Airy beam’s inherent nulls with co-channel users, achieving analog-domain interference suppression.
II-C2 Digital Precoding (Zero-Forcing)
Defining the effective channel as $\mathbf{H}_{\text{eff}}=\mathbf{H}_{\text{phys}}\mathbf{W}_{\text{RF}}$, we employ Zero-Forcing (ZF) to decouple user streams:
$\mathbf{W}_{\text{BB}}=\alpha\mathbf{H}_{\text{eff}}^{H}(\mathbf{H}_{\text{eff}}\mathbf{H}_{\text{eff}}^{H})^{-1},$ (5)
where $\alpha$ is the power normalization factor such that $\|\mathbf{W}_{\text{RF}}\mathbf{W}_{\text{BB}}\|_{F}^{2}=P_{\text{tx}}$.
II-D Performance Metrics and Analysis
To evaluate the system robustness under blockage, we utilize the following metrics.
II-D1 Condition Number ($\kappa$)
The condition number of the effective channel, $\kappa(\mathbf{H}_{\text{eff}})=\sigma_{\max}/\sigma_{\min}$, quantifies the orthogonality of the user subspaces. A high $\kappa$ (ill-conditioned channel) indicates that the beams intended for different users are highly correlated or that one user’s channel gain has collapsed due to blockage (singular value collapse).
II-D2 Signal-to-Interference-plus-Noise Ratio (SINR)
With the ZF precoding design described in (5), the effective channel is forced to be diagonal: $\mathbf{H}_{\text{eff}}\mathbf{W}_{\text{BB}}\approx\alpha\mathbf{I}_{K}$. Consequently, the received SINR for the $k$-th user is given by:
$\text{SINR}_{k}=\frac{|\alpha|^{2}}{\sigma^{2}}.$ (6)
Remark 2 (Equal SINR Implications):
The SINR expression in (6) corresponds to the post-precoding effective SINR under ZF baseband precoding. Specifically, with the ZF design in (5), the equivalent channel satisfies
$\mathbf{H}_{\mathrm{eff}}\mathbf{W}_{\mathrm{BB}}=\alpha\,\mathbf{I}_{K},$ (7)
which implies that all users experience the same effective gain $|\alpha|^{2}$ and hence identical SINR values. This does not imply that the underlying physical channels of shadowed and line-of-sight (LoS) users are identical. Instead, ZF enforces stream decoupling and equalization, while the channel disparity is absorbed into the global scaling factor $\alpha$. When $\mathbf{H}_{\mathrm{eff}}$ becomes ill-conditioned (e.g., under severe blockage), $\alpha$ sharply decreases due to noise amplification, leading to a simultaneous SINR degradation for all users.
We adopt this equalized ZF formulation to explicitly expose the effect of singular-value collapse and rank recovery in near-field multi-user channels. Alternative designs such as weighted ZF or explicit power allocation can yield unequal per-user SINR, but are outside the scope of this work.
III Baseline RNF Multiuser System with Conventional Focusing
This section establishes the performance baseline for a radiative near-field (RNF) multiuser (MU) system utilizing conventional focused beams in a blockage-free environment. By characterizing the system’s intrinsic spatial resolution using the rigorous Green’s function model (defined in Section II-B), we provide a theoretical upper bound. This benchmark is crucial for quantifying the severity of the blockage impairments discussed later in Section IV.
III-A Simulation Geometry and Parameters
We adopt the 2D simulation geometry illustrated in Fig. 2(a). The BS employs a large-scale ULA with $N=64$ elements and half-wavelength spacing ($d=0.49\lambda$), operating at $f_{c}=28$ GHz. This configuration yields a Fraunhofer distance of $R_{\text{FF}}\approx 1966\lambda$, placing the users deeply within the radiative near-field.
To ensure continuity with the blockage scenarios in subsequent sections, the user deployment is configured as follows:
  • UE-1 (Target User): Fixed at coordinates $\mathbf{p}_{1}=(-5\lambda,250\lambda)$. Remark: While this location is Line-of-Sight (LoS) in this baseline section, it is strategically chosen to fall within the “geometric shadow” region ($x<0$) once the half-space obstacle is introduced in Section IV.
  • UE-2 (Interfering User): Located at a different depth $z_{2}=300\lambda$. This user scans laterally along the $x$-axis from $x_{2}=-15\lambda$ (deep shadow side) to $10\lambda$ (bright side), crossing the angular line-of-sight of UE-1.
The system supports $N_{s}=2$ spatial streams with a normalized transmit power $P_{\text{tx}}=1$ and a noise floor $N_{0}$ set to $10^{-3}$ W, calibrated to provide a reference SNR of approximately 20 dB.
Figure 2: Baseline performance of the RNF-MU system with conventional focusing in free space. (a) System Geometry: UE-1 fixed at $(-5\lambda,250\lambda)$ and UE-2 scanning at $z=300\lambda$. (b) Condition Number $\kappa(\mathbf{H}_{\text{eff}})$: A sharp singularity ($\kappa_{\max}\approx 158$) occurs at $x_{2}=-6\lambda$, corresponding to angular alignment with UE-1. (c) Common SINR: Equalized SINR profile showing degradation at the angular collision. (d) Sum-Rate: Minimum throughput observed at the singularity point. (e)–(f) Angular Analysis: SINR and $\kappa$ plotted against UE-2’s angle $\theta_{2}$, confirming the overlap occurs at $\theta_{2}\approx\theta_{1}\approx-1.15^{\circ}$. (g)–(i) Field Intensity Distributions: Normalized field power $|E(x,z)|^{2}$ for UE-2 located at $x_{2}=-10\lambda$, $0\lambda$, and $10\lambda$, respectively, visualizing the distinct focal spots.
III-B Conventional Beamforming and Evaluation
III-B1 Green’s Function Channel Construction
In this blockage-free baseline, the physical channel vector $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ is generated directly using the free-space Green’s function model detailed in Section II-B. This ensures that the baseline evaluation captures the exact spherical wavefront curvature and path loss inherent to the RNF region, serving as a rigorous “ground truth” for comparison with the diffraction-based models in subsequent sections.
III-B2 Hybrid Precoding Strategy
The analog precoder $\mathbf{W}_{\text{RF}}$ employs Traditional Near-Field Focusing weights (phase conjugation) perfectly matched to each user’s location.
For the digital baseband precoder $\mathbf{W}_{\text{BB}}$, we employ Regularized Zero-Forcing (RZF) to mitigate multi-user interference while maintaining numerical stability near channel singularities. The unnormalized precoding matrix is given by:
$\tilde{\mathbf{W}}_{\text{BB}}=\mathbf{H}_{\text{eff}}^{H}\left(\mathbf{H}_{\text{eff}}\mathbf{H}_{\text{eff}}^{H}+\epsilon\mathbf{I}_{K}\right)^{-1},$ (8)
where $\epsilon\approx 10^{-10}$ is a regularization factor. The final precoder is power-normalized as $\mathbf{W}_{\text{BB}}=\alpha\tilde{\mathbf{W}}_{\text{BB}}$ to satisfy $\|\mathbf{W}_{\text{RF}}\mathbf{W}_{\text{BB}}\|_{F}^{2}=P_{\text{tx}}$.
III-C Performance Analysis: Resolution and Conditionality
Fig. 2 presents the comprehensive performance analysis.
III-C1 Condition Number and Angular Singularity
Fig. 2(b) plots the condition number $\kappa(\mathbf{H}_{\text{eff}})$ against the lateral position of UE-2. A critical observation is that the peak condition number ($\kappa\approx 158$) occurs at $x_{2}=-6\lambda$, not at $x_{2}=-5\lambda$. This shift is physically significant:
  • UE-1 is at $(-5\lambda,250\lambda)$, corresponding to an angle $\theta_{1}\approx\text{atan}(-5/250)\approx-1.15^{\circ}$.
  • The singularity occurs when UE-2 (at $z_{2}=300\lambda$) aligns with this angle: $x_{2}=300\lambda\cdot\tan(\theta_{1})=-6\lambda$.
Thus the collision occurs when UE-2 shares the same ray angle with UE-1, not when $x_{2}$ equals $x_{1}$. This indicates that for the given aperture ($D\approx 31\lambda$) and depth separation ($50\lambda$), traditional focusing relies primarily on angular separation to maintain orthogonality. The longitudinal resolution is insufficient to decorrelate users located on the same ray, leading to a “co-linear singularity.”
III-C2 Spectral Efficiency
Reflecting the conditioning analysis, the Sum-Rate (Fig. 2(d)) and Common SINR (Fig. 2(c)) exhibit a deep notch centered at the angular collision point ($x_{2}=-6\lambda$). Outside this narrow angular window (approx. $\pm 2^{\circ}$, as seen in Fig. 2(e)), the SINR recovers rapidly to its maximum plateau. This confirms that RNF MU-MIMO with traditional beams is highly effective but remains vulnerable to “flashlight effect” interference when users share the same Line-of-Sight angle.
Summary: The baseline analysis demonstrates that while traditional beams offer high gain, they suffer from rank collapse under angular alignment, even with RZF precoding. In the next section, we will see how a half-space obstacle exacerbates this by creating a “permanent” singularity zone (shadow) where traditional beams fail completely, motivating the need for Airy beamforming.
IV Half-Space Blockage: Double-Shadow Users and Airy-Geo Analog Beamforming
Having established the baseline performance in free space in Section III, we now investigate the critical scenario of half-space blockage. This section reveals the vulnerability of traditional near-field focusing when Line-of-Sight (LoS) paths are obstructed and demonstrates how the proposed Airy beamforming strategy—leveraging its auto-bending property—restores connectivity and recovers the channel rank in the geometric shadow region.
IV-A Half-Space Blockage Scenario and Singular Value Collapse
We consider a scenario where a semi-infinite obstacle (e.g., a wall or pillar) creates a half-space blockage. As illustrated in Fig. 3(a), a knife-edge obstacle is located at depth $z_{\text{obs}}=150\lambda$, blocking the entire region $x\leq x_{\text{obs}}=0$.
To rigorously evaluate the system’s resilience, we focus on a challenging “Double Shadow” configuration where both users are located deep within the geometric shadow zone. To ensure a fair comparison with the baseline in Section III:
  • UE-1 (Target): Remains fixed at $(x_{1},z_{1})=(-5\lambda,250\lambda)$, consistent with the baseline geometry (except now obstructed).
  • UE-2 (Scanning): Is located at depth $z_{2}=300\lambda$ but is now scanned laterally from $x_{2}=-15\lambda$ (deep shadow) to $x_{2}=-1\lambda$ (near the diffraction edge). This range is specifically chosen to investigate performance deep inside the shadow where traditional links fail.
In this setup, the direct LoS paths between the array and the users are geometrically severed. For traditional near-field focusing, which relies on linear propagation, energy transfer is restricted to weak diffraction components radiating from the obstacle’s edge. This leads to a catastrophic singular value collapse of the effective multi-user channel $\mathbf{H}_{\text{eff}}$. The channel matrix becomes ill-conditioned ($\kappa(\mathbf{H}_{\text{eff}})\to\infty$), and the ZF precoder fails to invert the channel without severely amplifying noise, resulting in a communication outage.
IV-B Proposed Airy-Geo Analog Beamforming Strategy
To overcome the shadow, we propose an “Airy-Geo” analog beamforming strategy. Unlike traditional beams that focus energy on a straight line, Airy beams follow a curved parabolic trajectory, enabling them to “ride” the diffraction edge and accelerate into the shadow zone.
IV-B1 Fixed Geometry-Based Codebook Design
A key feature of our proposal is the use of a simple, geometry-based analog codebook. We deliberately avoid complex, per-user optimization of the beam parameters to demonstrate the intrinsic robustness of the Airy wavepacket. The design follows two principles:
  1. 1.
    Physics-Informed Parameter Selection: We fix the curvature parameters $(B,F)$ for all users to maximize the “edge-riding” effect.
    • The focal length is set to $F\approx 163\lambda$ ($\approx$ 1.75 m at 28 GHz, $\lambda\approx$ 10.7 mm), slightly beyond the obstacle depth ($z_{\text{obs}}=150\lambda$). This ensures that the beam’s main lobe is fully formed and begins its transverse acceleration exactly as it passes the diffraction edge.
    • The bending coefficient is set to B=25B=-25. This relatively strong negative value imparts a significant transverse acceleration to the main lobe, bending it towards the positive $x$-direction (into the shadow) as it propagates.
  2. 2.
    Geometric Steering: For each user $k$, the initial launch angle $\theta_{k}$ is set simply to the geometric angle $\theta_{k}=\operatorname{atan2}(x_{k},z_{k})$, as plotted in Fig. 3(f). This confirms that our strategy relies purely on geometry without requiring complex channel state information for analog beam selection.
Remark 1 (Sub-optimality of Geometric Steering): It is important to acknowledge that in a blockage scenario, pointing the beam directly at the user ($\theta_{k}$) is not strictly optimal. The optimal strategy would likely involve aiming at a virtual source near the diffraction edge to maximize diffracted energy flux. However, we adopt this sub-optimal, geometry-based steering to highlight that even a constrained Airy design significantly outperforms traditional focusing due to its inherent self-healing physics.
Figure 3: Performance comparison in the double-shadow scenario ($x_{1}=-5\lambda$, fixed; $x_{2}$ scanning inside shadow). (a) System geometry showing the knife-edge blockage at $z=150\lambda$. (b) Sum-rate performance: Airy beams (red) maintain a robust rate floor in the deep shadow where traditional beams (blue) suffer outage. (c) Sum-rate gain of Airy over traditional focusing, showing gains up to 4 bps/Hz. (d) Common SINR: Airy beams provide a usable SINR floor despite interference, whereas traditional SINR collapses. (e) Condition number: Note that while Airy has a higher $\kappa$ near collision, it provides non-zero singular values, enabling communication. (f) Steering angles used for the Airy beams, following simple geometric logic.
Figure 4: Normalized field intensity distributions $|E(x,z)|^{2}$ (dB). Top Row (a, b): Deep shadow scenario ($x_{2}=-11\lambda$). Traditional beams (a) are severed, while Airy beams (b) bend significantly to serve the distant user. Bottom Row (c, d): Angular collision scenario ($x_{2}\approx x_{1}=-5\lambda$). Traditional beams (c) fail completely, while Airy beams (d) successfully deliver power to both co-located users despite the spatial overlap. Parameters: $B=-25,F\approx 163\lambda$.
IV-C Performance Evaluation and Analysis
We evaluate the performance using the rigorous Fresnel diffraction channel model described in Section II. To reflect a realistic low-SNR regime typical of non-LoS scenarios, the noise power is set to $N_{0}=10^{-3}$ (normalized). The results are summarized in Fig. 3 and visualized in Fig. 4.
IV-C1 SINR Resilience in Deep Shadow
Fig. 3(d) compares the Common SINR of the Traditional and Airy-Geo schemes.
  • Traditional Failure: As expected, the traditional focusing scheme (blue curve) experiences a severe outage in the deep shadow region ($x_{2}<-6\lambda$). The SINR drops significantly below 0 dB, indicating that the diffracted signal power is buried in the noise. This confirms that linear beams cannot effectively deliver energy behind the half-space blockage.
  • Airy Resilience: In contrast, the Airy scheme (red curve) maintains a usable SINR level (consistently above 0 dB) throughout the shadow region. This SINR resilience is critical: it implies that the Airy beam successfully establishes a viable communication link where traditional methods fail completely. While the absolute sum-rate (Fig. 3(b)) is naturally constrained by the high noise floor, the relative advantage of maintaining connectivity is significant.
IV-C2 Mechanism of Recovery: Active Bending
The physical mechanism behind this gain is visualized in Fig. 4.
  • Deep Shadow Recovery (Top Row): Fig. 4(a) and (b) correspond to a large separation ($x_{2}=-11\lambda$). Traditional beams (a) are sharply truncated, creating a dark triangular shadow zone. In contrast, Airy beams (b) exhibit a clear curved trajectory. The main lobe accelerates transversely, effectively “turning the corner” of the obstacle to illuminate the deep shadow user UE-2.
  • Recovery under Angular Collision (Bottom Row): Fig. 4(c) and (d) depict the scenario where UE-2 is close to UE-1 ($x_{2}\approx-5\lambda$). Here, both users are shadowed. While traditional beams (c) result in a complete double outage, the Airy beams (d) manage to illuminate both locations. Although the beams spatially overlap near the obstacle edge, they retain sufficient distinctness in the depth domain to allow the ZF precoder to separate the streams.
IV-C3 The Trade-off: Power vs. Orthogonality
In Fig. 3(e), the condition number $\kappa(\mathbf{H}_{\text{eff}})$ of the traditional near-field focusing scheme appears numerically smaller than that of the Airy-based scheme in the deep-shadow regime. This observation should be interpreted with care. For the traditional design, severe blockage drives both the largest and smallest singular values of $\mathbf{H}_{\text{eff}}$ close to zero, i.e., $\sigma_{\max}\approx 0$ and $\sigma_{\min}\approx 0$. As a result, the numerical value of $\kappa=\sigma_{\max}/\sigma_{\min}$ becomes highly sensitive to numerical truncation and noise-floor effects, and does not reflect a usable channel. Indeed, the corresponding absolute channel gain is extremely small, leading to SINR outage despite a seemingly moderate condition number.
In contrast, the Airy-based transmission significantly elevates $\sigma_{\min}$ above the noise floor by delivering non-negligible diffracted energy into the shadow region. While diffraction-induced beam spreading may increase channel correlation and hence yield a larger $\kappa$, the resulting effective channel is operational, supporting reliable communication with a non-zero SINR. Therefore, in deep-shadow scenarios, absolute channel gain and the recovery of the smallest singular value are more relevant performance indicators than the condition number alone. This distinction explains why the Airy-based scheme achieves superior sum-rate performance even when its condition number is numerically larger.
However, the SINR results in Fig. 3(d) conclusively demonstrate that recovering signal power is paramount in this noise-limited regime; the system achieves a viable link despite the increased interference, which is subsequently managed by the ZF precoder.
[…]
IV-C4 Summary
In summary, the proposed Airy-Geo strategy, even with a fixed and sub-optimal parameter configuration, successfully mitigates half-space blockage. By exploiting the auto-bending property, it transforms the “hard” geometric shadow into a “soft” coverage zone, providing a reliable fallback for RNF multi-user communications.
V Airy-Null Optimization in Mixed Shadow-Bright Scenarios
While Section IV demonstrated the efficacy of Airy beams in purely shadowed regions, a more complex challenge arises in mixed blockage scenarios, where some users are shadowed while others reside in the bright (LoS) region. In such cases, the extended oscillatory structure of the Airy beam—essential for its self-healing property—can cause severe interference to the LoS users. This section proposes an interference-aware optimization strategy that fine-tunes the Airy beam parameters to achieve analog-domain null steering.
V-A Scenario: The Interference Challenge
We consider a mixed geometry as depicted in Fig. 5(a):
  • UE-1 (Shadow): Located at $(-5\lambda,250\lambda)$, served by an Airy beam.
  • UE-2 (Bright - Hard Case): Located at $(+3.5\lambda,300\lambda)$. This position is termed a “hard case” because it lies close to the diffraction boundary, where the oscillating tail of UE-1’s Airy beam is strong.
  • Precoding Strategy: UE-2 uses a fixed traditional beam (optimal for LoS). The optimization focuses solely on UE-1’s Airy beam parameters.
If we apply the Airy-Geo strategy (geometric steering) from Section IV, the main lobe serves UE-1, but the side-lobes heavily pollute UE-2’s channel. This degrades the condition number of $\mathbf{H}_{\text{eff}}$, causing the ZF precoder to amplify noise ($N_{0}=10^{-3}$) and limiting the sum-rate.
V-B Problem Formulation: Joint Parameter Optimization
To mitigate this coupling, we exploit the parameterized structure of the Airy beam. The complex aperture field depends on the tuple $\bm{\psi}=(B,F,\theta_{1})$. Our goal is to find the optimal $\bm{\psi}^{*}$ that maximizes the sum-rate while maintaining a service guarantee for the shadowed user.
The optimization problem is formulated as:
$\max_{B,F,\theta_{1}}\quad$ $R_{\text{sum}}(\mathbf{H}_{\text{eff}}(B,F,\theta_{1}))$ (9a)
s.t. $|h_{11}(\bm{\psi})|^{2}\geq\eta\cdot|h_{11}^{\text{geo}}|^{2},$ (9b)
where $h_{11}^{\text{geo}}$ is the channel gain obtained using geometric steering, and $\eta\in[0.3,0.5]$ is a relaxation factor. This constraint ensures that the optimization does not trivially minimize interference by shutting down the target user’s link.
Input: Geometry; noise power $N_{0}$; transmit power $P_{\text{tx}}$; service constraint factor $\eta$; coarse grids $\mathcal{B}_{c},\mathcal{F}_{c},\Delta\Theta_{c}$; fine grids $\mathcal{B}_{f},\mathcal{F}_{f},\Delta\Theta_{f}$.
Output: Optimal UE-1 Airy parameters $(B^{\star},F^{\star},\theta_{1}^{\star})$.
1
2Step 0: Initialization & Baseline
3 $\theta_{1}\leftarrow\theta_{\text{geo}}$; compute baseline gain $|h_{11}^{\text{geo}}|^{2}$ and threshold $\tau\leftarrow\eta|h_{11}^{\text{geo}}|^{2}$
4 Fix UE-2 beam (traditional) and compute its channel column $\mathbf{h}_{2}$
5
6Step 1: Coarse Search (Global)
7 $R_{\text{best}}\leftarrow$-$\infty$
8 foreach $B\in\mathcal{B}_{c}$ do
9  foreach $F\in\mathcal{F}_{c}$ do
10     foreach $\Delta\theta\in\Delta\Theta_{c}$ do
11        Generate candidate UE-1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{geo}}+\Delta\theta)$
12        Propagate to obtain $\mathbf{h}_{1}$
13        if $|h_{11}|^{2}<\tau$ then
14          continue
15        end if
16       Form $\mathbf{H}_{\text{eff}}\leftarrow[\mathbf{h}_{1}\ \mathbf{h}_{2}]$ and compute sum-rate $R$
17        if $R>R_{\text{best}}$ then
18           Update $(B_{\text{best}},F_{\text{best}},\theta_{\text{best}})\leftarrow(B,F,\theta_{\text{geo}}+\Delta\theta)$
19           $R_{\text{best}}\leftarrow R$
20
21        end if
22
23     end foreach
24
25  end foreach
26
27 end foreach
28
29Step 2: Fine Search (Local Refinement)
30 Define fine grids around $(B_{\text{best}},F_{\text{best}},\theta_{\text{best}})$ to form $\mathcal{B}_{f},\mathcal{F}_{f},\Delta\Theta_{f}$
31 foreach $B\in\mathcal{B}_{f}$ do
32  foreach $F\in\mathcal{F}_{f}$ do
33     foreach $\Delta\theta\in\Delta\Theta_{f}$ do
34        Generate candidate UE-1 Airy beam $E_{1}(x)$ with parameters $(B,F,\theta_{\text{best}}+\Delta\theta)$
35        Propagate to obtain $\mathbf{h}_{1}$
36        if $|h_{11}|^{2}<\tau$ then
37          continue
38        end if
39       Form $\mathbf{H}_{\text{eff}}\leftarrow[\mathbf{h}_{1}\ \mathbf{h}_{2}]$ and compute sum-rate $R$
40        if $R>R_{\text{best}}$ then
41           Update $(B_{\text{best}},F_{\text{best}},\theta_{\text{best}})\leftarrow(B,F,\theta_{\text{best}}+\Delta\theta)$
42           $R_{\text{best}}\leftarrow R$
43
44        end if
45
46     end foreach
47
48  end foreach
49
50 end foreach
51
52Return:
53 Set $(B^{\star},F^{\star},\theta_{1}^{\star})\leftarrow(B_{\text{best}},F_{\text{best}},\theta_{\text{best}})$
54
Algorithm 1 Coarse-to-Fine Airy-Null Optimization
V-C Proposed Solution: Coarse-to-Fine Search
Since the objective function is non-convex and oscillatory due to the phase structure of Airy beams, gradient-based methods may plunge into local optima. To address this, we propose a robust two-stage search strategy, summarized in Algorithm 1.
V-D Complexity Analysis
The computational efficiency of the proposed beamforming strategy is a critical factor for practical implementation. Unlike fully digital precoding schemes that require optimizing complex-valued weights for every antenna element, the proposed coarse-to-fine Airy-null search optimizes only three scalar analog parameters $(B,F,\theta_{1})$.
Let $N_{\text{coarse}}=|\mathcal{B}_{c}||\mathcal{F}_{c}||\Delta\Theta_{c}|$ and $N_{\text{fine}}=|\mathcal{B}_{f}||\mathcal{F}_{f}||\Delta\Theta_{f}|$ denote the number of candidate parameter tuples evaluated in the coarse and fine search stages, respectively, consistent with the sets defined in Algorithm 1. For each candidate tuple, the dominant computational cost lies in the evaluation of the Fresnel propagation operator $\mathcal{P}_{z}$ (see Eq. (3)). Implementing this operator via the FFT-based Angular Spectrum Method on a transverse aperture grid of size $N_{x}$ (where $N_{x}\gg N$ is chosen to ensure sufficient sampling density and mitigate boundary reflections) yields a propagation complexity of $\mathcal{O}(N_{x}\log N_{x})$. Consequently, the total computational complexity of the proposed algorithm scales as:
$C_{\text{total}}=\mathcal{O}\left((N_{\text{coarse}}+N_{\text{fine}})N_{x}\log N_{x}\right).$ (10)
It is worth noting that since the traditional focused beam for the Line-of-Sight (LoS) user (UE-2) is fixed during the optimization of the shadowed user (UE-1), its effective channel column $\mathbf{h}_{2}$ can be precomputed once. This reduces the per-iteration cost solely to the update of the Airy beam column $\mathbf{h}_{1}$.
Comparison with Conventional Schemes: In contrast, a conventional fully digital multi-user beamforming optimization typically involves $N\times K$ complex variables. Solving for the optimal precoder often requires iterative algorithms (e.g., Weighted MMSE or SCA) involving matrix inversions, which scale with a complexity of at least $\mathcal{O}(KN^{3})$ or $\mathcal{O}(K^{2}N^{2})$ per iteration. For ELAA systems where $N$ is large (e.g., $N\geq 256$), such operations are impractical for real-time. Our proposed method, leveraging the parameterized physics of Airy beams, drastically reduces the dimensionality of the search space. Furthermore, the optimization can be performed offline to generate a geometry-based codebook (Lookup Table), rendering the online complexity negligible ($\mathcal{O}(1)$) for static blockage scenarios.
Figure 5: Airy-Null optimization results in the mixed Shadow-Bright scenario. (a) Geometry showing UE-1 in shadow and UE-2 in bright region. (b) Sum-rate vs. steering angle $\Delta\theta$: The peak occurs at $\Delta\theta\approx-2.8^{\circ}$, not at the geometric zero. (c) Condition number $\kappa$ vs. $\theta$: A sharp dip in $\kappa$ coincides with the sum-rate peak, indicating improved orthogonality. (d)-(e) Comparison of Sum-Rate and Condition Number for Trad-All, Airy-Geo, and Airy-Opt strategies. (f) Analog coupling strengths, highlighting the interference suppression capability of the optimized design.
V-E Results and Physical Insight: Energy Rebalancing
The performance of the proposed Airy-Opt strategy is compared against the Trad-All (baseline) and Airy-Geo (unoptimized) schemes in Fig. 5 and Fig. 6. The results reveal a fundamental mechanism of sum-rate maximization that goes beyond simple interference nulling.
V-E1 Sum-Rate Gain and Interference Suppression
Fig. 5(d) shows that Airy-Opt significantly outperforms both Trad-All (which fails due to blockage) and Airy-Geo (which suffers from interference). A key observation from Fig. 5(f) is the analog coupling strength: the optimized design effectively suppresses the interference leakage $|h_{21}|^{2}$ from the shadowed UE-1 to the bright UE-2.
V-E2 Physical Mechanism: Service Rebalancing
To understand the root cause of this gain, we analyze the transverse field intensity distribution at the depth of UE-2 ($z=300\lambda$), as visualized in Fig. 6. The comparison between Airy-Geo (gray curve) and Airy-Opt (red curve) reveals a crucial energy rebalancing effect:
  • At UE-1 (Shadow, $x=-5\lambda$): The Airy-Geo beam delivers very weak energy (approx. 32-32 dB) because its trajectory is not optimally bent for this specific blockage geometry. In contrast, the Airy-Opt beam significantly boosts the signal power to 11-11 dB, a massive gain of +21+21 dB. This confirms that the optimization successfully “pulls” the main lobe energy back into the deep shadow region.
  • At UE-2 (Bright, $x=+3.5\lambda$): The Airy-Geo beam creates a strong side-lobe peak (8.6-8.6 dB) at the location of the bright user. The Airy-Opt beam, by adjusting its curvature and angle, reduces this interference to 12.3-12.3 dB. While this is a reduction of only $3.7$ dB, it is sufficient to mitigate interference without sacrificing the primary link.
Core Insight: The essence of the optimization is not merely to place a perfect deep null at UE-2, but to rebalance the service between the two users. The geometric solution (Airy-Geo) is overly biased towards the bright region (due to the diffraction nature), starving the shadowed user. The optimized solution (Airy-Opt) sacrifices a small amount of bright-region energy (which is abundant) to critically resurrect the shadowed link, thereby maximizing the overall system throughput.
Figure 6: Comparison of normalized transverse field intensity at the depth of UE-2 ($z=300\lambda$). The curves show the beam profiles for the geometric solution (Airy-Geo, gray) and the optimized solution (Airy-Opt, red). The optimization effectively “pulls” energy back into the shadow region to boost UE-1 (+21+21 dB) while slightly reducing the interference level at the bright-region user UE-2 (3.7-3.7 dB), thereby rebalancing the service quality.
Figure 7: Robustness analysis against lateral positioning errors of the bright-region user (UE-2). The beamformer design assumes a nominal position $\hat{\mathbf{p}}_{2}$, while the true position varies by $\Delta x_{2}$. (a) Sum-rate $R_{\text{sum}}$ versus positioning error $\Delta x_{2}/\lambda$: Airy-Opt maintains a robust rate floor across the entire error range. (b) Condition number $\kappa(\mathbf{H}_{\text{eff}})$ versus $\Delta x_{2}/\lambda$: Airy-Opt exhibits a consistently low and flat $\kappa$, indicating stable channel separability. (c) Sum-rate gain relative to the traditional baseline: Airy-Opt provides a sustained gain of >3.5>3.5\,bps/Hz. (d) Worst-case sum-rate (over $\pm\Delta x_{2}$) versus error magnitude: The optimized design significantly outperforms both geometric and traditional benchmarks even under large estimation mismatches.
V-F Robustness to User Positioning Errors
The proposed Airy-null optimization strategy (Section V-C) relies on the spatial distinctness of the users to shape the interference in the analog domain. In practical RNF systems, however, channel state information (CSI) is inevitably imperfect. A critical question arises: Does the fine-tuned Airy beam become hypersensitive to positioning errors, causing the performance to collapse if the user is slightly misaligned? To address this, we evaluate the system’s robustness against lateral positioning errors of the bright-region user (UE-2).
V-F1 Error Model and Evaluation Setup
We consider a scenario where the beamformer design is based on an estimated position $\hat{\mathbf{p}}_{2}=(x_{2},z_{2})$, while the actual user is located at $\mathbf{p}_{2}=(x_{2}+\Delta x_{2},z_{2})$. We sweep the lateral error $\Delta x_{2}$ within the range $3\lambda,+3\lambda]$. The setup assumes the “hard case” geometry defined in Section V-A, where the nominal positions are fixed, and the shadowed user (UE-1) is assumed to be accurately tracked to isolate the impact of interference leakage.
Three strategies are compared under identical error conditions:
  • Trad-All: Both users are served by traditional near-field focused beams designed for their estimated positions.
  • Airy-Geo: UE-1 employs the geometric Airy beam (B=25,F=1.75B=-25,F=1.75\,m, $\theta=\theta_{\text{geo}}$) while UE-2 uses the traditional beam.
  • Airy-Opt: UE-1 employs the optimized Airy beam (B=44,F=1.50B=-44,F=1.50\,m, $\theta=\theta_{\text{geo}}-2.9^{\circ}$) derived from Algorithm 1, while UE-2 uses the traditional beam.
Crucially, the precoder for UE-2 ($w_{2}$) is fixed to the estimated position $\hat{\mathbf{p}}_{2}$ for all schemes, meaning it does not adapt to the error, thereby creating a realistic mismatch scenario.
V-F2 Results and Physical Interpretation
Fig. 7 presents the impact of positioning errors on the sum-rate and channel conditioning.
Sum-Rate Resilience: As shown in Fig. 7(a), the Airy-Opt strategy maintains a superior sum-rate across the entire error sweep range $|\Delta x_{2}|\leq 3\lambda$. While the geometric solution (Airy-Geo) degrades rapidly as the user moves into the Airy beam’s strong side-lobe region (positive $\Delta x$), the optimized beam exhibits remarkable stability. Even in the worst-case error scenario (Fig. 7(d)), Airy-Opt retains a sum-rate advantage of approximately 3344 bps/Hz over the traditional baseline.
Channel Separability: The mechanism behind this resilience is elucidated in Fig. 7(b), which plots the condition number $\kappa(\mathbf{H}_{\text{eff}})$ of the effective channel. The traditional scheme suffers from severe ill-conditioning ($\kappa>200$) regardless of small displacements, as the fundamental rank deficiency persists. The Airy-Geo scheme shows a rising $\kappa$ as the error increases, indicating that interference leakage is creating collinearity. In contrast, Airy-Opt maintains a consistently low and flat condition number profile ($\kappa\approx 3$). This indicates that the optimization does not merely find a “pinhole” null; rather, by jointly tuning $(B,F,\theta)$, it reshapes the wavefront to create a broader “soft” interference-suppression zone.
Gain Retention: Fig. 7(c) illustrates the sum-rate gain of the Airy strategies relative to the traditional baseline. The optimized Airy beam delivers a stable gain (mean +3.89+3.89 bps/Hz) that is largely insensitive to the specific direction of the error. This confirms that the proposed method achieves analog-domain interference shaping, effectively rebalancing the power-interference trade-off in a robust manner. By “pulling” the energy into the shadow (for UE-1) while smoothing the side-lobe fluctuations in the bright region (near UE-2), the Airy-Opt design ensures that the multi-user cluster remains separable even under realistic estimation uncertainties.
VI Conclusion and Future Work
In this paper, we have investigated the fundamental limitations of traditional radiative near-field (RNF) focusing in half-space blockage scenarios and proposed a novel Airy-beam-assisted transmission strategy to overcome these challenges. Our analysis, grounded in rigorous Fresnel diffraction modeling, revealed that conventional focused beams suffer from a catastrophic “singular value collapse” when encountering knife-edge obstacles, rendering zero-forcing precoding ineffective for shadowed users.
To address this, we exploited the unique auto-bending and self-healing properties of Airy beams. In the double-shadow scenario (Section IV), we demonstrated that a geometry-based Airy beamforming strategy can effectively “ride” the obstacle edge, delivering significant signal power into the geometric shadow and restoring the rank of the multi-user channel. Furthermore, in the more complex mixed shadow-bright scenario (Section V), we proposed an interference-aware optimization framework. Our results showed that by jointly optimizing the curvature and launch angle, the system can achieve a critical energy rebalancing—sacrificing a negligible amount of power in the bright region to achieve massive gains (over 20 dB) for the shadowed user, thereby maximizing the system sum-rate.
This work establishes Airy beams as a promising, hardware-efficient alternative to RIS for blockage mitigation in future 6G near-field networks. Future research directions include extending this framework to 3D blockage geometries, investigating the integration of Airy beams with dynamic user tracking, and exploring learning-based methods for real-time beam parameter adaptation in rapidly changing environments.